OpenStudy (kanwal32):

if the quadratic equation ax^2+bx+6=0 does not have two distinct real roots, then the least value of 2a+b is

OpenStudy (kanwal32):

@ganeshie8

OpenStudy (kanwal32):

@nincompoop @Luigi0210

OpenStudy (kanwal32):

@dan815

OpenStudy (kanwal32):

@UnkleRhaukus

OpenStudy (unklerhaukus):

what is the discriminant of that quadratic? \[\Delta = b^2-4ac\]

OpenStudy (kanwal32):

b^2-24a<0

OpenStudy (unklerhaukus):

when ∆ > 0 the are two real roots there is only one root when ∆ = 0 is ∆ < 0 there are no real roots

OpenStudy (kanwal32):

yes b^2-4ac<0

OpenStudy (kanwal32):

then

OpenStudy (unklerhaukus):

i think you want \[b^2-24a\leq0\]

OpenStudy (kanwal32):

yes so how to find 2a+b least value

OpenStudy (unklerhaukus):

solve for a (as a function of b)

OpenStudy (kanwal32):

how can we find

OpenStudy (unklerhaukus):

solve the inequality

OpenStudy (kanwal32):

can u solve for this

OpenStudy (kanwal32):

\[-2\sqrt{6}a<b<2\sqrt{6}a\]

OpenStudy (kanwal32):

how???????????

OpenStudy (unklerhaukus):

\[b^2-24a\leq0\\ b^2\leq24a\\ \frac{b^2}{24}\leq a\]

OpenStudy (kanwal32):

then

OpenStudy (unklerhaukus):

2a + b =

OpenStudy (unklerhaukus):

\[2a+b\geq2\cdot\frac{b^2}{12}+b\]

OpenStudy (unklerhaukus):

following?, not following? @kanwal32